OpenStudy (anonymous):

lim sin4x/sin3x x-->0

OpenStudy (campbell_st):

find a substitute for sin(4x) say 2sin(2x)cos(2x) and sin(3x) say sin(2x + x) = sin(2x)cos(x) + cos(2x)sin(x) maybe expanding further and cancelling... to simplify...

OpenStudy (campbell_st):

here is something that may help http://mathworld.wolfram.com/Multiple-AngleFormulas.html

OpenStudy (dumbcow):

L'Hopitals may work as well differentiate top and bottom \[\rightarrow \frac{4\cos(4x)}{3\cos(3x)}\] now take limit